Explain how to prepare 800 ml of 0.78 mmol/litre NAD+ (MW 655.43) solution

1
Expert's answer
2022-01-28T01:07:02-0500

According to the equation:


M = m / (Mr × V)


where M - molarity, m - mass, Mr - molecular weight, V - volume.


From here:


m = M × Mr × V




1. m(NAD+) = 0.78 mmol/l × 655.43 g/mol × 800 ml = 0.78 × 10-3 mol/l × 655.43 g/mol × 0.8 l = 0.41 g


To prepare 800 ml of 0.78 mmol/ll NAD+ solution, it is required to take 0.41 g of NAD and adjust water up to 800 ml.


2. m(ATP) = 0.55 mmol/l × 555.14 g/mol × 100 ml = 0.55 × 10-3 mol/l × 555.14 g/mol × 0.1 l = 0.03 g


To prepare 100 ml of 0.55 mmol/litre ATP solution, it is required to take 0.03 g of ATP and adjust water up to 100 ml.


3. m(Tris HCl) = 100 mmol/l × 157.6 g/mol × 400 ml = 6.3 g


m(Mg acetate) = 1 mmol/ × 214.45 g/mol × 400 ml = 0.086 g


m(NAD+) = 0.78 mmol/l × 655.43 g/mol × 400 ml = 0.205 g


m(ATP) = 0.55 mmol/l × 555.14 g/mo × 400 ml = 0.122 g


To prepare 400 ml of the buffer solution, it is required to take 6.3 g of Tris HCl, 0.086 g of magnesium acetate, 0.205 g of NAD, 0.122 g of ATP and adjust with water up to 400 ml.